The time,t, in seconds that it takes a body to fall a distance of d meters from rest is given by the equation t=(2d/g)^1/2,where g is the acceleration due to gravity in m/s^2. Find d when t = 6 and g = 9.8. Explain how you arrived at your answer
1
Expert's answer
2012-02-02T11:08:48-0500
The time, t, in seconds that it takes a body to fall a distance of d meters from rest is given by the equation t=(2d/g)^1/2,where g is the acceleration due to gravity in m/s^2. Find d when t = 6 and g = 9.8.

We just need to apply the given formula:

t = (2d/g)^1/2 ==> t² = 2d/g ==> d = gt²/2 = 9.8*6²/2 = 352.8 m.